# Theorem of the heart for Weibel’s homotopy -theory

### 09 Mar 2026 v1

@@ Abstract.
The proof is heavily based on another new result, which is a much stronger version of Barwick’s theorem of the heart. Its special case states the following: if is a small stable category with a bounded -structure, such that for some the realization functor induces isomorphisms on between the objects of then the map is an isomorphism for and a monomorphism for Moreover, we prove that these estimates are sharp, even for dg categories over a field. In particular, the naive -theoretic theorem of the heart fails for
@@ 1 Coherently assembled abelian categories and exact categories
- 1.4 Criterion of -accessibility for exact categories
@@ 0. Introduction
Originally, Weibel’s definition of was given in the -linear context, see [Wei89, Wei13, Tab15]. The natural generalization for stable categories is explained for example in [E25c, Section 8, Corollary 8.3]. For a small stable category the spectrum is given by the geometric realization
@@
We note that the above results are new only when the hearts are non-noetherian (and non-artinian). Namely, if is noetherian, then we have by [AGH19, Theorem 3.6], and we have (more precisely, for all ).
We prove that the estimates in Corollaries 0.4 and 0.5 are sharp (Theorem 7.1), even if we are dealing with small dg categories resp. small abelian categories over a field. In particular, we obtain that the naive -theoretic theorem of the heart fails for which is a much more precise version of a result of Ramzi-Sosnilo-Winges [RSW24, Theorem B].
We now explain a very surprising analogy with the familiar results on connective -ring spectra. Intuitively the notion of a -structure is dual to the notion of a weight structure [Bon10, Sos19]. In particular, the bounded -structures are “dual” to bounded weight structures. Consider an idempotent-complete stable category with a bounded weight structure For simplicity we assume that is generated by a single object (this assumption is harmless, it will make no essential difference in what follows). In this case we may assume that is of weight zero. Putting (meaning the spectrum of endomorphisms), we obtain and is a connective -ring. Conversely, for a connective -ring the category has a unique bounded weight structure such that has weight zero [Sos19].
@@
It is clear from the formulations that our Theorem 0.3 for some is “dual” to Proposition 0.6 for both in the assumptions and in the conclusion one simply replaces with for the relevant spectra, and one replaces epimorphisms with monomorphisms. Continuing this analogy, Corollary 0.4 for some is “dual” to the special case of Proposition 0.6 for the same when and for Further, Corollary 0.5 is “dual” to the case and We do not have any conceptual explanation for why such a duality takes place, but we note that in both contexts the (co)connectivity estimates are sharp.
Now we explain why from this point of view our Theorem 0.1 is “dual” to the Dundas-Goodwillie-McCarthy theorem. We recall some notation. We denote by the category of idempotent-complete stable categories and consider the universal finitary localizing invariant [BGT13, E25c]. The category is naturally symmetric monoidal, and moreover it is rigid (in the sense of Gaitsgory and Rozenblyum [GaiRoz17, Definition 9.12]) by [E25c, Theorem 0.1]. We also consider the universal finitary -invariant localizing invariant so is a quotient of by the cocomplete ideal generated by the reduced motive of the flat affine line By [E25c, Theorem 8.1] this is a smashing localization. More precisely we have the idempotent -algebra in given by the geometric realization where We have an equivalence
@@
where and is the topological cyclic homology [BHM93, NS18]. The map is the cyclotomic trace. In view of the isomorphism (0.2), the DGM theorem states that we have an equivalence
@@
We mention the following natural questions which are not studied in this paper. The first one is somewhat vague, and it is motivated by the paper [LS25] and by the above analogy (duality) with weight structures.
@@
The following question is motivated by the above analogy with connective -rings. We introduce some non-standard notation. Let be a (not necessarily exact) functor between -presentable stable categories, commuting with -filtered colimits. For denote by the universal polynomial functor of degree commuting with -filtered colimits, with a map (its existence follows from the adjoint functor theorem). We say that the coanalytic approximation of is the colimit (computed objectwise).
@@
In Section 1 we introduce and study locally -coherent exact -categories, focusing on coherently assembled exact -categories. We are mostly interested in the presentable stable envelope of a coherently assembled exact -category and in the properties of this category, such as dualizability. In particular, we define the continuous -theory of such exact -categories by putting Our approach in this section is somewhat minimalistic, in particular we mostly avoid the non-coherently assembled case. We reduce most of the non-trivial statements to the familiar results about small exact -categories and their (small) stable envelopes. Among other things, we explain a natural version of the situation of dévissage for coherently assembled abelian categories in Subsection 1.9. The results of this section are used in the proofs of our main theorems, especially in Section 3.
@@ 1.2. Accessibly exact categories
We define a locally -coherent exact category to be a category of the form where is a small exact category and is a regular cardinal. The exact structure is described in Proposition 1.9.
@@
We prove (i). If is -accessible stable with standard exact structure, then the statement is evident. The general case reduces to the stable case by the proof of Proposition 1.9: we have a fully faithful -accessible functor and This proves (i).
@@
In the situation of Proposition 1.15 we will consider as a directed union of over all regular
@@ 1.3. Grothendieck exact categories
Any presentable stable category can be considered as a Grothendieck exact category. More generally, any Grothendieck prestable category [Lur18, Definition C.1.4.2] is a Grothendieck exact category.
1.4. Criterion of -accessibility for exact categories
@@ 1.5. Presentable stable envelopes
Let us temporarily denote by the right hand side of (1.3). Choose a regular cardinal such that is locally -coherent exact. Then restriction to defines a fully faithful functor
@@
Recall that for a Grothendeick abelian category the so-called unseparated derived category can be described as a (dg nerve of) the dg category of complexes of injective objects of It has a natural right complete -structure, compatible with filtered colimits, and we have Moreover, is generated by the heart as a localizing subcategory. We refer to [Kr05, Kr15] and [Lur18, Section C.5.8] for a detailed account.
@@
Essentially the same argument as in loc. cit. proves a similar result for the so-called Grothendieck abelian -categories, i.e. categories of the form where is a prestable Grothendieck category and [Lur18, Definition C.5.4.1]. Namely, if we equip with the induced exact structure from then is equivalent to where is left adjoint to Here resp. is the category of Grothendieck prestable categories resp. Grothendieck abelian -categories, where the -morphisms are colimit-preserving left exact functors.
@@ 1.7. Coherently assembled exact categories
Suppose that is locally -coherent. Consider the retraction Both functors commute with filtered colimits. By Remark 1.29 the functor is exact, and by definition of a Grothendieck exact category the functor is exact. This proves the implication.
Since and are continuous, the functor is isomorphic to the composition
@@
Now consider the general case. Choose a retraction where is additive and compactly generated, and are additive and continuous, and is moreover strongly continuous (e.g. take ). Denote by the category of cofiltered limit-preserving functors As in Proposition 1.23 we see that we have a universal continuous additive functor Consider also a similar functor The latter is simply the ind-completion of the functor where is equipped with the split exact structure. It follows from the retraction that is fully faithful and its essential image is closed under extensions. Consider the induced exact structure on Using the retraction again we see that is a coherently assembled exact category, and any short exact sequence in is a filtered colimit of split short exact sequences. It remains to obtain the stated description of exact projections in
@@
Conversely, left be a morphism in such that any compact morphism factors through Let where is directed. Each map is compact, hence so is by the strong continuity of Moreover, for each the map factors through It follows from the above special case that is an exact projection in and applying we conclude that is an exact projection in ∎
@@
The category from the proof of Proposition 1.32 has a natural -structure such that the connective part is identified with the category of cofiltered limit-preserving functors Moreover, is a dualizable module over i.e. a dualizable object of – the symmetric monoidal category of additive presentable -categories [LLS]. Moreover, any dualizable additive category can be obtained in this way: one recovers as a full subcategory of flat objects of Here an object is flat if the functor commutes with finite limits.
@@
(ii) Put We need to show that for we have an isomorphism It suffices to consider the case where and is the universal exact functor. Then the isomorphism follows directly from (i): we have a commutative square of compactly assembled categories
@@
From now on, for a coherently assembled exact category we consider the small category as an exact category with the induced exact structure from and similarly for any regular Note that by Proposition 1.15 for any uncountable regular cardinals the functor is fully faithful, and we consider as a directed union of
@@
Next, if is a -small direct sum of some then it suffices to take to be the direct sum of where the projection is as above. Now, consider the general case, i.e. where is a -small directed poset, and It suffices to prove that the map
@@
Let be a coherently assembled exact category, and let be a presentable stable category. Suppose that we have a continuous exact functor which is fully faithful, reflects exactness and its essential image is closed under extensions. Consider the induced (continuous exact) functor and denote by the universal continuous exact functor, so that . Then for the map
@@ 1.8. Coherently assembled abelian categories
is coherently assembled and for any object the map is a monomorphism in
@@
. Since is locally noetherian, it is locally coherent, hence coherently assembled by Proposition 1.43. Any object is a directed union of noetherian subobjects hence the map is a monomorphism.
. The assumption implies that the functor
@@
Now take some and let where is directed. Since the map is a monomorphism, we may and will assume that each map is a monomorphism in For each the map is zero, hence by the left exactness of Therefore, each is compact and is compactly generated. By Proposition 1.43, is locally coherent. It remains to show that the class of compact objects in is closed under subobjects. This again follows from the left exactness of if is compact and then hence i.e. is compact. ∎
@@
Note that is a (strictly) full subcategory. The following result is an (almost) straightforward generalization of the following fact: if is an ind-system of (ordinary) associative coherent rings with left flat transition maps, then the ring is right coherent.
@@ 1.9. The situation of dévissage
For any consider the full subcategory consisting of pairs such that Then we have a direct sequence in By Proposition 1.49, the functor is fully faithful. It is an equivalence since is generated by via filtered colimits. By Proposition 1.48 it suffices to prove that is coherently assembled for each
@@ 2.1. Reminder on -structures
Recall that for a -structure on the left completion is given by the limit This is a stable category, and it naturally has a -structure where We have a natural exact and -exact functor If it is an equivalence, then the -structure on is called left complete. Dually, we have a notion of a right completion and similarly for right completeness. Note that in general the right completion of is equivalent to the left completion of and both are given by
@@
Let be a presentable stable category, and let be an accessible -structure compatible with filtered colimits. We denote by the localizing subcategory generated by and similarly for
@@
Next, 2.3 follows directly from 2.1. Namely, the inclusion is colimit-preserving, hence it has a right adjoint. The right orthogonal to its image is identified with or equivalently with It follows from 2.1 that the image of in is contained in the right orthogonal to Since and generate we obtain the semi-orthogonal decomposition (2.2). The equivalence (2.3) follows. ∎
@@ 2.3. Compactly assembled -structures
We need to show that for we have This follows from the isomorphisms in
@@ 2.4. General constructions of -structures
It remains to prove that for any there exists some and a functor such that and we have for where If this is the case for some object we say that has a good filtration. Since generates as a stable subcategory, we only need to show that for any object with a good filtration, for any for any and for any morphism the object has a good filtration.
@@
By adjunction, the left -exactness of the functor implies that the composition is fully faithful. Since its image generates the target as a localizing subcategory, the assertion is a special case of Proposition 2.22. ∎
@@ 3. Refined -theoretic theorem of the heart for dualizable categories
Let be a dualizable category with a compactly assembled continuously bounded -structure Let be an integer. Then the map is an isomorphism for and a monomorphism for
@@
Let be a dualizable -category. Suppose that for some we have isomorphisms for and (this assumption always holds for ) Then the map is an isomorphism for and a monomorphism for
@@
Our arguments are completely different from [Bar15]. Even if we are only interested in still the the proof of [Bar15, Theorem 6.1] does not generalize to dualizable (not compactly generated) categories. We also note that an example below (Corollary 8.2) shows that the group might be non-zero for a dualizable -category. In other words, the statement of [AGH19, Theorem 2.35] does not hold for dualizable categories. However, a different version of a construction from loc. cit. will be very useful for us, see Propositions 3.4 and 3.5 below.
@@
Let be as in Theorem 3.1. Consider the small stable category with the bounded -structure induced from and let be its simultaneous left and right completion, i.e. . Equivalently, is the left completion of the category i.e.
@@
Let be as in Theorem 3.1 and let be an integer. Consider the exact category whose underlying category is and a morphism is an inclusion in if and only if both morphisms and are monomorphisms in
@@
It is clear that is a strongly continuous exact functor (by definition of the exact category ). Hence, by Proposition 1.34 induces a strongly continuous exact functor By Proposition 2.22 it remains to prove that the induced maps are isomorphisms for and (where can be strictly negative). We take such and denote by the universal functor. By 3.5 we have a cartesian square of spectra:
@@
Next, we prove that for the map is an isomorphism for each and a monomorphism for By Proposition 3.7 the category from (3.7) has a compactly assembled continuously bounded s-structure, and we have an equivalence By the assertion on from loc. cit. and by Proposition 2.19 we have an equivalence for Hence, by Lemma 3.3 we have a surjection for On the other hand, the composition is left inverse to the realization functor. Therefore the map from (3.7) induces an isomorphism for The long exact sequence of homotopy groups for (3.7) shows that the map is an isomorphism for and a monomorphism for as stated.

### 16 Mar 2026 v2

@@ Abstract.
The proof is heavily based on another new result, which is a much stronger version of Barwick’s theorem of the heart. Its special case states the following: if is a small stable category with a bounded -structure, such that for some the realization functor induces isomorphisms on between the objects of then the map is an isomorphism for and a monomorphism for Moreover, we prove that these estimates are sharp, even for dg categories over a field. In particular the naive -theoretic theorem of the heart fails for
@@ 1 Coherently assembled abelian categories and exact categories
- 1.4 Criterion for an exact category to be locally -coherent
@@ 0. Introduction
Originally, Weibel’s definition of was given in the -linear context [Wei89, Wei13, Tab15]. The natural generalization for stable categories is explained for example in [E25c, Section 8, Corollary 8.3]. For a small stable category the spectrum is given by the geometric realization
@@
We note that the above results are new only when the hearts are non-noetherian (and non-artinian). Namely, if is noetherian, then we have by a theorem of Antieau-Gepner-Heller [AGH19, Theorem 3.6], and we have (more precisely, for all ).
We prove that the estimates in Corollaries 0.4 and 0.5 are sharp (Theorem 7.1), even if we are dealing with small dg categories resp. small abelian categories over a field. In particular we obtain that the naive -theoretic theorem of the heart fails for which is a much more precise version of a result of Ramzi-Sosnilo-Winges [RSW24, Theorem B].
We now explain a very surprising analogy with the familiar results on connective -ring spectra. Intuitively the notion of a -structure is dual to the notion of a weight structure [Bon10, Sos19]. In particular the bounded -structures are “dual” to bounded weight structures. Consider an idempotent-complete stable category with a bounded weight structure For simplicity we assume that is generated by a single object (this assumption is harmless, it will make no essential difference in what follows). In this case we may assume that is of weight zero. Putting (meaning the spectrum of endomorphisms), we obtain and is a connective -ring. Conversely, for a connective -ring the category has a unique bounded weight structure such that has weight zero [Sos19].
@@
It is clear from the formulations that our Theorem 0.3 for some is “dual” to Proposition 0.6 for both in the assumptions and in the conclusions one simply replaces with for the relevant spectra, and one replaces epimorphisms with monomorphisms. Continuing this analogy, Corollary 0.4 for some is “dual” to the special case of Proposition 0.6 for the same when and for Further, Corollary 0.5 is “dual” to the case and We do not have any conceptual explanation for why such a duality takes place, but we note that in both contexts the (co)connectivity estimates are sharp.
Now we explain why from this point of view our Theorem 0.1 is “dual” to the Dundas-Goodwillie-McCarthy theorem. We recall some notation. We denote by the category of idempotent-complete stable categories and consider the universal finitary localizing invariant [BGT13, E25c]. The category is naturally symmetric monoidal, and moreover it is rigid by [E25c, Theorem 0.1] (we mean rigidity in the sense of Gaitsgory and Rozenblyum [GaiRoz17, Definition 9.12]). We also consider the universal finitary -invariant localizing invariant so is a quotient of by the cocomplete ideal generated by the reduced motive of the flat affine line By [E25c, Theorem 8.1] this is a smashing localization. More precisely, we have the idempotent -algebra in given by the geometric realization where We have an equivalence
@@
where and is the topological cyclic homology [BHM93, NS18]. The map is the cyclotomic trace. In view of the equivalence (0.2), the DGM theorem states that we have an equivalence
@@
We mention the following natural questions which are not studied in this paper. The first one is somewhat vague, and it is motivated by the work of Levy and Sosnilo [LS25] and by the above analogy (duality) with weight structures.
@@
The following question is motivated by the above analogy with connective -rings and by the functor (topological restriction) with coefficients in a bimodule [LM12, KMN23]. We introduce some non-standard notation for the “dual” of the familiar Goodwillie tower [Goo03]. Let be a (not necessarily exact) functor between -presentable stable categories, commuting with -filtered colimits. For denote by the universal polynomial functor of degree commuting with -filtered colimits, with a map (its existence follows from the adjoint functor theorem). We say that the coanalytic approximation of is the colimit (computed objectwise).
@@
In Section 1 we introduce and study locally -coherent exact -categories, focusing on coherently assembled exact -categories. We are mostly interested in the presentable stable envelope when is coherently assembled, and in the properties of the category such as dualizability. In particular, we define the continuous -theory of such exact -categories by putting Our approach in this section is somewhat minimalistic, in particular we mostly avoid the non-coherently assembled case. We reduce most of the non-trivial statements to the familiar results about small exact -categories and their (small) stable envelopes. Among other things, we explain a natural version of the situation of dévissage for coherently assembled abelian categories in Subsection 1.9. The results of this section are used in the proofs of our main theorems, especially in Section 3.
@@ 1.2. Accessibly exact categories
We define a locally -coherent exact category to be a category of the form where is a small exact category and is a regular cardinal. The exact structure is described in Proposition 1.9. For we say “locally coherent” instead of “locally -coherent”.
@@
We prove (i). If is -accessible stable with the standard exact structure, then the statement is evident. The general case reduces to the stable case by the proof of Proposition 1.9: we have a fully faithful -accessible functor and This proves (i).
@@
In the situation of Proposition 1.15 we define the category to be the directed union of over all regular
@@ 1.3. Grothendieck exact categories
Any presentable stable category can be considered as a Grothendieck exact category. More generally, any Grothendieck prestable category ([Lur18, Definition C.1.4.2]) is a Grothendieck exact category.
1.4. Criterion for an exact category to be locally -coherent
@@ 1.5. Presentable stable envelopes
Let us temporarily denote by the right hand side of (1.3). Choose a regular cardinal such that is locally -coherent. Then restriction to defines a fully faithful functor
@@
Recall that for a Grothendeick abelian category the so-called unseparated derived category can be described as (the dg nerve of) the dg category of complexes of injective objects of It has a natural right complete -structure, compatible with filtered colimits, and we have Moreover, is generated by the heart as a localizing subcategory. We refer to [Kr05, Kr15] and [Lur18, Section C.5.8] for a detailed account.
@@
Essentially the same argument as in loc. cit. proves a similar result for the so-called Grothendieck abelian -categories, i.e. categories of the form where is a Grothendieck prestable category and [Lur18, Definition C.5.4.1]. Namely, if we equip with the induced exact structure from then is equivalent to where is left adjoint to Here resp. is the category of Grothendieck prestable categories resp. Grothendieck abelian -categories, where the -morphisms are colimit-preserving left exact functors.
@@ 1.7. Coherently assembled exact categories
. Suppose that is locally -coherent. Consider the retraction Both functors commute with filtered colimits. By Remark 1.29 the functor is exact, and by definition of a Grothendieck exact category the functor is exact. This proves the implication.
. Since and are continuous, the functor is isomorphic to the composition
@@
Now consider the general case. Choose a retraction where is additive and compactly generated, and are additive and continuous, and is moreover strongly continuous (e.g. take ). Denote by the category of cofiltered limit-preserving functors As in Proposition 1.23 we see that we have a universal continuous additive functor Consider also a similar functor The latter is simply the ind-completion of the functor where is equipped with the split exact structure. It follows from the retraction that is fully faithful and its essential image is closed under extensions. Consider the induced exact structure on Using the retraction again we see that is a coherently assembled exact category (by Proposition 1.31), and any short exact sequence in is a filtered colimit of split short exact sequences. It remains to obtain the stated description of exact projections in
@@
Conversely, left be a morphism in such that any compact morphism factors through Let where is directed. Each map is compact, hence so is by the strong continuity of Moreover, for each the map factors through It follows from the above special case that is an exact projection in Applying we conclude that is an exact projection in ∎
@@
The category from the proof of Proposition 1.32 has a natural -structure such that the connective part is identified with the category of cofiltered limit-preserving functors Moreover, is a dualizable module over i.e. a dualizable object of – the symmetric monoidal category of additive presentable -categories [LLS]. Moreover, any dualizable additive category can be obtained in this way: one recovers as the full subcategory of flat objects of Here an object is flat if the functor commutes with finite limits.
@@
(ii) Put We need to show that for we have an isomorphism It suffices to consider the case where and is the universal continuous exact functor. Then the isomorphism follows directly from (i): we have a commutative square of compactly assembled categories
@@
From now on, for a coherently assembled exact category we consider the small category as an exact category with the induced exact structure from and similarly for any regular Recall that by Proposition 1.15 for any uncountable regular cardinals the functor is fully faithful, and we defined to be the directed union of
@@
Next, if is a -small direct sum of some then it suffices to take to be the direct sum of where the projections are as above. Now consider the general case, i.e. where is a -small directed poset, and It suffices to prove that the map
@@
Let be a coherently assembled exact category, and let be a presentable stable category. Suppose that we have a continuous exact functor which is fully faithful, reflects exactness and its essential image is closed under extensions. Consider the induced (continuous exact) functor and denote by the universal continuous exact functor, so that Then for the map
@@ 1.8. Coherently assembled abelian categories
is compactly assembled and for any object the map is a monomorphism in
@@
. Any object is a directed union of noetherian subobjects hence the map is a monomorphism.
. We first observe that is coherently assembled. Indeed, the functor is automatically right exact, so we only need to check that it preserves monomorphisms. If is a monomorphism in then in the commutative square
both vertical arrows and the lower horizontal arrow are monomorphisms, hence so is the upper horizontal arrow, as required.
Next, we see that the functor
@@
Now take some and let where is directed. Since the map is a monomorphism, we may and will assume that each map is a monomorphism in For each the map is zero, hence by the left exactness of Therefore, each is compact and is compactly generated. By Proposition 1.43 is locally coherent. It remains to show that the class of compact objects in is closed under subobjects. This again follows from the left exactness of if is compact and then hence i.e. is compact. ∎
@@
Note that is a (strictly) full subcategory. The following result is an (almost) straightforward generalization of the following fact: if is an ind-system of (ordinary) associative right coherent rings with left flat transition maps, then the ring is right coherent.
@@ 1.9. The situation of dévissage
For any consider the full subcategory consisting of pairs such that Then we have a direct sequence in By Proposition 1.49 the functor is fully faithful. It is an equivalence since is generated by all via filtered colimits. By Proposition 1.48 it suffices to prove that is coherently assembled for each
@@ 2.1. Reminder on -structures
Recall that for a -structure on the left completion is given by the limit This is a stable category, and it naturally has a -structure where We have a natural exact and -exact functor If it is an equivalence, then the -structure on is called left complete. Dually, we have a notion of the right completion and similarly for right completeness. Note that in general the right completion of is equivalent to the left completion of and both are given by
@@
Let be a presentable stable category, and let be an accessible -structure compatible with filtered colimits. We denote by the localizing subcategory generated by and similarly for and
@@
Next, 2.3 follows directly from 2.1. Namely, the inclusion is colimit-preserving, hence it has a right adjoint. The right orthogonal is identified with or equivalently with It follows from 2.1 that the image of in is contained in the right orthogonal to Since and generate we obtain the semi-orthogonal decomposition (2.2). The equivalence (2.3) follows. ∎
@@ 2.3. Compactly assembled -structures
The “if” direction follows from Proposition 2.6. For the “only if” direction we need to show that for we have This follows from the isomorphisms in
@@ 2.4. General constructions of -structures
As usual we denote by the poset with the usual order. We only need to prove that for any there exists some and a functor such that and we have for where If this is the case for some object we say that has a good filtration. Since generates as a stable subcategory, we only need to show that for any object with a good filtration, for any for any and for any morphism the object has a good filtration.
@@
By adjunction, the left -exactness of the functor implies that the composition is fully faithful. Since its image generates the target as a localizing subcategory, the assertion of the corollary is a special case of Proposition 2.22. ∎
@@ 3. Refined -theoretic theorem of the heart for dualizable categories
Let be a dualizable -category. Let be an integer. Then the map is an isomorphism for and a monomorphism for
@@
Let be a dualizable -category. Suppose that for some we have isomorphisms for and (this assumption always holds for ). Then the map is an isomorphism for and a monomorphism for
@@
Our arguments are completely different from [Bar15]. Even if we are only interested in still the proof of [Bar15, Theorem 6.1] does not generalize to dualizable (not compactly generated) categories. We also note that an example below (Corollary 8.2) shows that the group might be non-zero for a dualizable -category. In other words, the statement of [AGH19, Theorem 2.35] does not hold for dualizable categories. However, a different version of a construction from loc. cit. will be very useful for us, see Propositions 3.4 and 3.5 below.
@@
Let be a dualizable -category. Consider the small stable category with the bounded -structure induced from and let be its simultaneous left and right completion, i.e. . Equivalently, is the left completion of the category i.e.
@@
Let be a dualizable -category and let be an integer. Consider the exact category whose underlying category is and the exact structure is the following. A morphism is an inclusion in if and only if both morphisms and are monomorphisms in Further, is a projection in if and only if both morphisms and are epimorphisms in
@@
It is clear that is a strongly continuous exact functor (by definition of the exact category ). Hence, by Propositions 1.26 and 1.34 induces a strongly continuous exact functor By Proposition 2.22 it remains to prove that the induced maps are isomorphisms for and (where can be strictly negative). We take such and denote by the universal functor. By 3.5 we have a cartesian square of spectra:
@@
Next, we prove that for the map is an isomorphism for each and a monomorphism for By Proposition 3.7 the category from (3.7) has a compactly assembled continuously bounded -structure, and we have an equivalence By the assertion on from loc. cit. and by Proposition 2.19 we have an equivalence for Hence, by Lemma 3.3 we have a surjection for On the other hand, the composition is left inverse to the realization functor. Therefore the map from (3.7) induces an isomorphism for The long exact sequence of homotopy groups for (3.7) shows that the map is an isomorphism for and a monomorphism for as stated.